# Lifshitz tail states in non-Hermitian disordered photonic lattices

### v1

@@ Abstract
In recent years, non-Hermitian systems –characterized by complex potentials– have emerged as a fertile ground for novel wave phenomena, driven by their ability to model energy gain, loss, and nonreciprocal effects [21]. Non-Hermitian extensions of Anderson localization have uncovered phenomena distinct from their Hermitian counterparts, including the non-Hermitian delocalization transition [22, 23, 24, 25] and a form of universal non-Hermitian transport [26, 27, 28, 29, 30]. In these systems, wave propagation can occur in a jumpy manner even when the eigenstates of the Hamiltonian are exponentially localized [26, 30]. Lifshitz tail states are also present in non-Hermitian disordered lattices [31, 32]. Although their physical significance and practical relevance have remained uncertain [32], a recent study [30] highlighted the role of the density of state tails in establishing the dynamical scaling in non-Hermitian transport.
In this Letter, we show that Lifshitz tail states, despite being extreme events, can dominate the dynamical behavior of disordered non-Hermitian lattices of finite size. Specifically, we focus on Anderson localization induced by imaginary on-site potential Bernoulli disorder in non-Hermitian photonic lattices. We demonstrate that, in such systems, Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the overall system behavior. These findings challenge conventional views of Lifshitz tail states as negligible and highlight their experimental observability in non-Hermitian settings. To illustrate this phenomenon, we analyze the Anderson-Bernoulli model, a paradigmatic framework for studying localization in a two-component alloy [18, 20, 33, 34], adapted here to a non-Hermitian photonic lattice. Our results provide insights into the dynamics of non-Hermitian systems with disorder and reveal potential avenues for harnessing non-Hermitian models to visualize rare events like Lifshitz tail states.
@@
where $\psi_{n}(t)$ is the wave amplitude at the $n$-th lattice site ($n=1,2,..,L$), $J$ is the hopping rate, $V_{n}$ is the on-site random potential, and $\psi_{0}(t)=\psi_{L+1}(t)=0$ for open boundary conditions. $V_{n}$ are assumed to be independent random variables with the same probability density distribution. For the Bernoulli-Anderson model [20, 33, 34], $V_{n}$ can take only two values, $V_{n}=W$ with probability $p$ and $V_{n}=0$ with probability $(1-p)$. Hence the on-site potential is a stochastic sequence of 0 and $W$. In the non-Hermitian Anderson-Bernoulli model, the potential is imaginary, namely $W=-i\gamma$ where $\gamma>0$ is the particle loss rate at site $n$. In the photonic context, this model basically describes a coherent random walk of a photon on a lattice with traps [35], which can be experimentally implemented in synthetic temporal mesh lattices (see e.g. [15, 26, 36, 37] and reference therein).
The eigenvalues $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ are complex with ${\rm Im}(\lambda)\leq 0$, and we are interested in the behavior of $\lambda$ and corresponding eigenvectors $u_{n}$ in the large system size ($L\rightarrow\infty$) limit. For a large loss rate $\gamma\gg J$, the eigenvalues $\lambda$ and corresponding eigenvectors $u_{n}$ form two distinct groups, one with ${\rm Im}(\lambda)\simeq-\gamma$ corresponding to short-lived states, and the other one with ${\rm Im}(\lambda)\simeq 0^{-}$ corresponding to long-lived states; see Fig.1(a). The eigenvectors of the former group have their excitations mainly in the lossy sites, where $V_{n}=-i\gamma$, while the latter group of eigenstates have their excitations mostly in the lossless sites, where $V_{n}=0$. The energy spectrum of both groups is symmetric about the real axis ${\rm Re}(\lambda)=0$. In fact, it can be readily shown that, if $u_{n}$ is an eigenfunction of $\mathcal{H}$ with eigenvalue $\lambda$, then $(-1)^{n}u^{*}_{n}$ is an eigenvector of $\mathcal{H}$ with eigenvalue $-\lambda^{*}$.
@@
Figure 1: (a) Behavior of the energy spectrum $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ in complex plane for a given realization of the random Bernoulli potential $V_{n}$. Parameter values are $\gamma/J=5$, $p=0.2$ and $L=5000$. The spectrum is symmetry around the imaginary axis ${\rm{Re}}(\lambda)=0$ and can be grouped in two sets: the eigenvalues with small imaginary parts ($|{\rm Im}(\lambda)|\sim J^{2}/\gamma$), corresponding to short-lived eigenstates, and the eigenvalues with large imaginary parts ($|{\rm Im}(\lambda)|\sim\gamma$), corresponding to long-lived eigenstates. The Lifshitz tail eigenstates are those with eigenvalues $\lambda$ close to the upper left and right corners, possessing extremely long lifetimes. The insets in (a) shows an enlargement of the spectrum near the upper right corner. The dashed curve in the panel depicts the loci of the eigenenergies of the Lifshitz tail states predicted by the "particle-in-a-box" model [Eq.(2)]. (b) Numerically-computed behavior of the Lifshitz tail state possessing the longest lifetime (modulus of $|u_{n}|$, red circles). The position of the traps is also shown for comparison. This eigenstate is well approximated by the ground state of a potential well of size $l_{m}$ (dashed rectangular curve), corresponding to the longest sequence of consecutive zeros in the disorder realization $V_{n}$ (blue squares).
@@
System dynamics and Lifshitz tail delocalization. The delocalization features of the Lifshitz tail states are hard to be observed in large-size Hermitian systems, since they are extremely rare states and, for a given realization of the disorder sequence $V_{n}$, their spatial locations in the lattice are unknown, rendering their excitation unfeasible. Conversely, in non-Hermitian lattices Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the long-time system behavior. In fact, let us assume that at initial time all sites of the lattice are equally excited, i.e. let us assume $\psi_{n}(0)=(1/\sqrt{L})$. This condition ensures that Lifshitz tail states of the right edge band are surely excited, regardless of their unknown location and size in the lattice. The solution to the Schrödinger equation (1) reads
@@
i.e. the long-time system’s behavior is dominated by the most rare Lifshitz tail state. The emergence of such a state from the system’s dynamics can be detected by monitoring the temporal behavior of the dynamical IPR, which is defined as
@@
Figure 2: (a) Numerically-computed behavior of the IPR of the Lifshitz tail state with the longest lifetime versus lattice size $L$ (circles); the IPR is averaged over 200 realizations of Bernoulli disorder. Parameter values are $p=0.2$ and $\gamma/J=5$. The solid curve depicts the theoretically predicted behavior based on Eqs.(3) and (4). (b) Dynamical evolution of $|\psi_{n}(t)|$ in a lattice of size $L=500$ for a single realization of Bernoulli disorder. Initial condition is $\psi_{n}(0)=1/\sqrt{L}$. The upper panel in (b) shows the behavior of $|\psi_{n}|$ at final propagation time $t/J=1200$ (red circles), and the realization of Bernoulli disorder (blue squares). The main peak corresponds to the ground state of the potential box with largest size $l=l_{m}$ of the disorder realization. (c) Numerically-computed behavior of the dynamical IPR (solid curve), averaged over 100 realizations of disorder. The dashed horizontal curve in the asymptotic theoretical value predicted by Eq.(8).
@@
In the above equations, $u_{n}^{(m)}$ and $w_{n}^{(m)}$ are the pulse amplitudes at discrete time step $m$ and lattice site (unit cell) $n$ in the two fiber loops, $\beta$ is the coupling angle and $\gamma_{n}$ are uncorrelated stochastic losses (photonic traps) taking the two values $\gamma_{n}=\gamma>0$ with probability $p$ and $\gamma_{n}=0$ with probability $(1-p)$. A finite lattice, containing $L$ unit cells, is obtained by assuming a coupling angle $\beta=\pi/2$ at the lattice edges, so that in Eqs.(9) and (10) the site index $n$ can be restricted to the range $n=1,2,...,L$. Like for the continuous random walk with traps [Eq.(1)], the long-time dynamics of the discrete random walk defined by Eqs.(9) and (10) is governed by the eigenstate of the propagator $\mathcal{U}=\exp(-i\mathcal{H})$ with the longest lifetime, which is localized at the potential box of the Bernoulli distribution with the largest size $l=l_{m}$. An example of such an eigenstate, for a given realization of the trap distribution in the lattice, is shown in Fig.3(a). Such a state can be observed in the long-time random walk dynamics when one of the two fiber loops is seeded with a long flat-top optical pulse, corresponding to the initial condition $u_{n}^{(0)}=1/\sqrt{L}$ and $w_{n}^{(0)}=0$, as shown in Fig.3(b). In fact, the asymptotic temporal dynamics is dominated by the eigenstate of $\mathcal{U}$ with the longest lifetime, which is localized in the potential box of largest size $l_{m}$, as illustrated in Fig.3(a). Figure 3(c) shows the numerically-computed time evolution of the dynamical IPR, averaged over 200 realizations of disorder, clearly indicating stabilization to an asymptotic value which is in good agreement with the theoretical prediction [Eq.(8)].
Figure 3: (a) Numerically-computed behavior of the eigenstate (red circles) of the propagator $\mathcal{U}$ with the longest lifetime for the discrete-time photonic random walk with traps. The position of the traps in the lattice are denoted by the blue squares. Parameter values are $\beta=\pi/6$, $L=100$, $p=0.2$ and $\gamma=2$. (b) Temporal light dynamics in the synthetic lattice (behavior of $I_{n}=|u_{n}^{(m)}|^{2}+|w_{n}^{(m)}|^{2}$, normalized to its norm, versus propagation time step $m$ on a pseudocolor map) for uniform initial excitation $u_{n}^{(0)}=1/\sqrt{L}$, $w_{n}^{(0)}=0$. (c) Temporal evolution of the dynamical IPR versus time step $m$, averaged over 200 realizations of disorder. The IPR is defined as ${\rm IPR}=\sum_{n}I_{n}^{2}/\sum_{n}I_{n}$. The dashed horizontal curve is the asymptotic value predicted by Eq.(8).

### v2

@@ Abstract
In recent years, non-Hermitian systems –characterized by complex potentials– have emerged as a fertile ground for novel wave phenomena, driven by their ability to model energy gain, loss, and nonreciprocal effects [21]. Non-Hermitian extensions of Anderson localization have uncovered phenomena distinct from their Hermitian counterparts, including the non-Hermitian delocalization transition [22, 23, 24, 25] and a form of universal non-Hermitian transport [26, 27, 28, 29, 30, 31]. In these systems, wave propagation can occur in a jumpy manner even when the eigenstates of the Hamiltonian are exponentially localized [26, 30]. Lifshitz tail states are also present in non-Hermitian disordered lattices [32, 33]. Although their physical significance and practical relevance have remained uncertain [33], recent studies [30, 31] highlighted the role of the density of state tails in establishing the dynamical scaling in non-Hermitian transport.
In this Letter, we show that Lifshitz tail states, despite being extreme events, can dominate the
dynamical behavior of disordered non-Hermitian lattices of finite size. Specifically, we focus on Anderson localization induced by imaginary on-site potential Bernoulli disorder in non-Hermitian photonic lattices. We demonstrate that, in such systems, Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the overall system behavior. These findings challenge conventional views of Lifshitz tail states as negligible and highlight their experimental observability in non-Hermitian settings. To illustrate this phenomenon, we analyze the Anderson-Bernoulli model, a paradigmatic framework for studying localization in a two-component alloy [18, 20, 34, 35], adapted here to a non-Hermitian photonic lattice. Our results provide insights into the dynamics of non-Hermitian systems with disorder and reveal potential avenues for harnessing non-Hermitian models to visualize rare events like Lifshitz tail states.
@@
where $\psi_{n}(t)$ is the wave amplitude at the $n$-th lattice site ($n=1,2,..,L$), $J$ is the hopping rate, $V_{n}$ is the on-site random potential, and $\psi_{0}(t)=\psi_{L+1}(t)=0$ for open boundary conditions. $V_{n}$ are assumed to be independent random variables with the same probability density distribution. For the Bernoulli-Anderson model [20, 34, 35], $V_{n}$ can take only two values, $V_{n}=W$ with probability $p$ and $V_{n}=0$ with probability $(1-p)$. Hence the on-site potential is a stochastic sequence of 0 and $W$. In the dissipative Anderson-Bernoulli model, the potential is imaginary, i.e. $W=-i\gamma$ where $\gamma>0$ is the particle loss rate. In this case, a photon occupying a site with $V_{n}=-i\gamma$ has a non-vanishing probability of being annihilated, thus providing a trap for photons [36]. Such a trapped configuration can be experimentally implemented in optical waveguide arrays with tailored losses or in dissipative synthetic lattices, such as in temporal mesh lattices [15, 26, 37, 38]. The eigenvalues $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ are complex with ${\rm Im}(\lambda)\leq 0$, and we are interested in the behavior of $\lambda$ and corresponding eigenvectors $u_{n}$ in the large system size ($L\rightarrow\infty$) limit. For a large loss rate $\gamma\gg J$, the eigenvalues $\lambda$ and corresponding eigenvectors $u_{n}$ form two distinct groups, one with ${\rm Im}(\lambda)\simeq-\gamma$ corresponding to short-lived states, and the other one with ${\rm Im}(\lambda)\simeq 0^{-}$ corresponding to long-lived states; see Fig.1(a). The eigenvectors of the former group have their excitations mainly in the lossy sites, where $V_{n}=-i\gamma$, while the latter group of eigenstates have their excitations mostly in the lossless sites, where $V_{n}=0$. The energy spectrum of both groups is symmetric about the imaginary axis ${\rm Re}(\lambda)=0$, which follows from the symmetry $V_{n}=-V_{n}^{*}$. As shown in the Supplemental document, this symmetry is however not relevant for the formation of Lifshitz tail states.
@@
Figure 1: (a) Behavior of the energy spectrum $\lambda$ of the matrix Hamiltonian $\mathcal{H}$ in complex plane for a given realization of the random Bernoulli potential $V_{n}$. Parameter values are $\gamma/J=5$, $p=0.2$ and $L=5000$. The spectrum is symmetry around the imaginary axis ${\rm{Re}}(\lambda)=0$ and can be grouped in two sets: the eigenvalues with small imaginary parts ($|{\rm Im}(\lambda)|\sim J^{2}/\gamma$), corresponding to short-lived eigenstates, and the eigenvalues with large imaginary parts ($|{\rm Im}(\lambda)|\sim\gamma$), corresponding to long-lived eigenstates. The Lifshitz tail eigenstates are those with eigenvalues $\lambda$ close to the upper left and right corners, possessing extremely long lifetimes. The insets in (a) shows an enlargement of the spectrum near the upper right corner. The dashed curve in the panel depicts the loci of the eigenenergies of the Lifshitz tail states predicted by the "particle-in-a-box" model [Eq.(2)]. (b) Numerically-computed behavior of the Lifshitz tail state possessing the longest lifetime (modulus of $|u_{n}|$, red circles). The position of the traps is shown by filled blue rectangles. This eigenstate is well approximated by the ground state of a potential well of size $l_{m}$ (dashed rectangular curve), corresponding to the longest sequence with missing traps.
@@
System dynamics and Lifshitz tail delocalization. The delocalization features of the Lifshitz tail states are hard to be observed in large-size Hermitian systems, since they are extremely rare states and, for a given realization of the sequence $V_{n}$, their spatial locations in the lattice are unknown, rendering their excitation unfeasible. Conversely, in non-Hermitian lattices Lifshitz tail states not only become more prominent but also play a pivotal role in shaping the long-time system behavior. In fact, let us assume that at initial time all sites of the lattice are equally excited, i.e. let us assume $\psi_{n}(0)=(1/\sqrt{L})$. This condition ensures that Lifshitz tail states of the right edge band are surely excited, regardless of their unknown location and size in the lattice. The solution to the Schrödinger equation (1) reads
@@
i.e. the long-time system’s behavior is dominated by the most rare Lifshitz tail state. We note that, since the decay rate of Lifshitz states condense to zero as $L\rightarrow\infty$, a longer propagation time will be required to reach the steady state as the system size $L$ is increased. The emergence of such a state from the system’s dynamics is a general phenomenon, observable for $W$ larger than $\sim J$, and can be detected by monitoring the temporal behavior of the dynamical IPR, which is defined as
@@
Figure 2: (a) Numerically-computed behavior of the IPR of the Lifshitz tail state with the longest lifetime versus lattice size $L$ (circles), from $L=100$ to $L=2100$; the IPR is averaged over 200 realizations of Bernoulli disorder. Parameter values are $p=0.2$ and $\gamma/J=5$. The solid curve depicts the theoretically predicted behavior [Eqs.(3) and (4)]. (b) Dynamical evolution of $|\psi_{n}(t)|$ in a lattice of size $L=500$ for a single realization of Bernoulli disorder. Initial condition is $\psi_{n}(0)=1/\sqrt{L}$. The upper panel in (b) shows the behavior of $|\psi_{n}|$ at final propagation time $t=1200/J$ (red circles), and position of traps (blue squares). The main peak corresponds to the ground state of the potential box with largest size $l=l_{m}$. (c) Numerically-computed behavior of the dynamical IPR (solid curve), averaged over 100 realizations of disorder. The dashed horizontal curve in the asymptotic theoretical value predicted by Eq.(8).
@@
In the above analysis wee assumed a Bernoulli distribution of disorder, however Lifshitz tail states could be observed for other types of disorder, such as for a uniform imaginary disorder (see Supplemental document).
@@
In the above equations, $u_{n}^{(m)}$ and $w_{n}^{(m)}$ are the pulse amplitudes at discrete time step $m$ and lattice site (unit cell) $n$ in the two fiber loops, $\beta$ is the coupling angle and $\gamma_{n}$ are uncorrelated stochastic losses (photonic traps) taking the two values $\gamma_{n}=\gamma>0$ with probability $p$ and $\gamma_{n}=0$ with probability $(1-p)$. A finite lattice, containing $L$ unit cells, is obtained by assuming a coupling angle $\beta\neq\pi/2$ in the bulk, and $\beta=\pi/2$ at the two lattice edges. Like for the continuous random walk with traps [Eq.(1)], the long-time dynamics of the discrete random walk defined by Eqs.(9) and (10) is governed by the eigenstate of the propagator $\mathcal{U}=\exp(-i\mathcal{H})$ with the longest lifetime, which is localized at the potential box of the Bernoulli distribution with the largest size $l=l_{m}$. An example of such an eigenstate is shown in Fig.3(a). This state is observed in the long-time random walk dynamics when one of the two fiber loops is seeded with a long optical pulse, such a flat-top or a Gaussian pulse, as shown in Fig.3(b) for the initial uniform condition $u_{n}^{(0)}=1/\sqrt{L}$, $w_{n}^{(0)}=0$ and in Fig.S4 of the Supplemental document for an input Gaussian excitation. Figure 3(c) shows the numerically-computed time evolution of the dynamical IPR, averaged over 200 realizations of disorder, clearly indicating stabilization to an asymptotic value which is in good agreement with the theoretical prediction [Eq.(8)].
Figure 3: (a) Numerically-computed behavior of the eigenstate (red circles) of the propagator $\mathcal{U}$ with the longest lifetime for the dissipative photonic random walk. The position of the traps in the lattice are denoted by the blue squares. Parameter values are $\beta=\pi/6$, $L=100$, $p=0.2$ and $\gamma=2$. (b) Temporal light dynamics (behavior of $I_{n}=|u_{n}^{(m)}|^{2}+|w_{n}^{(m)}|^{2}$, normalized to its norm, versus propagation time step $m$ on a pseudocolor map) for uniform initial excitation $u_{n}^{(0)}=1/\sqrt{L}$, $w_{n}^{(0)}=0$. (c) Temporal evolution of the dynamical IPR versus time step $m$, averaged over 200 realizations of disorder. The IPR is defined as ${\rm IPR}=\sum_{n}I_{n}^{2}/\sum_{n}I_{n}$. The dashed horizontal curve is the asymptotic value predicted by Eq.(8).